# A greedy heuristic for graph burning

### v1

@@ 4.2 A greedy heuristic for GBP
It is well known that $b(P_{n})=\lceil n^{1/2}\rceil$ for any path $P_{n}$ of order $n$ [1]. So, given a path $P_{n}$, we can run Gr with $k=b(P_{n})$, which is equivalent to execute Algorithm 4 with $l=h=b(P_{n})$. For the rest of the proof, let us consider paths $P_{n}$ with the largest possible order for any given burning number, namely, paths $P_{n}$ with $n=z^{2}$ for any $z\in\mathbb{Z}^{+}$. Also, observe that the maximum number of vertices a vertex $v$ can cover in a path is $|N_{r}[v]|=2r+1$, for any given $r$. The proof is divided into three parts. The first considers the initial iteration of the for loop of Gr ($r=k-1$); the second considers iterations $r=k-2$ to $r=2$, and the last one considers iterations $r=1$ and $r=0$. In the first iteration, $r=k-1$, Gr selects a vertex that covers exactly $12r+$ 1 vertices because the path’s length, $(r+1)^{2}$, is bounded by $12r+$ 1, i.e.,
@@
At iterations $r=k-2$ to $r=2$, Gr selects a vertex with maximum coverage from the largest subpath induced by removing the covered vertices. The largest subpath has at least $12r+$ 1 vertices, i.e.,
@@
where $(r+2)^{2}$ is the number of uncovered vertices two iterations before, and $12(r+1)+$ 1 is the number of newly covered vertices at the previous iteration. Thus, the numerator is the number of uncovered vertices at the beginning of iteration $r$. This way, at iteration $r$, Gr selects a vertex that covers $12r+$ 1 vertices, none of which were covered in the previous iterations. Notice that ties can be broken at random. By mathematical induction, Eq. (20) holds for any integer $r\geq 2$. At iteration $r=1$, the number of previously covered vertices equals $\sum_{r=2}^{k-1}(2r+1)=k^{2}-4$. Namely, at this point, only four vertices are left uncovered. By direct inspection, there will be cases where only two or three new vertices can be covered at iterations $r=1$ and $r=0$, leaving the rest uncovered (See Figure 3). Therefore, the constructed solution might not be a burning sequence.
Figure 3: A non-burning sequence returned by Gr over $P_{25}$ with $k=b(P_{25})=5$. At the end, two vertices are uncovered.
@@
Since a cycle $C_{n}$ is a path $P_{n}$ with its endpoints joined, it follows that $b(C_{n})=\lceil n^{1/2}\rceil$ [1]. So, we can run Gr with $k=b(C_{n})$, which is equivalent to execute Algorithm 4 with $l=h=b(C_{n})$. The first vertex selected by Gr maximizes the coverage. Then, removing the covered vertices induces a path $P$. By Proposition 5, Gr might return a non-burning sequence of length $b(C_{n})-1$ over $P$, which is concatenated to the first selected vertex, becoming a non-burning sequence of length $b(C_{n})$ for $C_{n}$. ∎
By Proposition 5 and Corollary 1, Gr might not seem very appealing. However, as reported in Section 5, it performed unexpectedly well over benchmark and synthetic graphs. Anyhow, we can do better by repeating Gr up to $n$ times and selecting a different initial vertex every time; we refer to this heuristic as GrP (Algorithm 6). Once the first burning sequence of length $k$ is found (line 3), exploring the remaining vertices is unnecessary. Nevertheless, the complexity of GrP embedded into a binary search (Algorithm 4) is $\mathcal{O}(n^{4}\log n)$, which is impractical for sufficiently large graphs.
@@
1:A graph $G=(V,E)$, its distance matrix $d$, and a guess $k$ on $b(G)$
@@
4:     $s\leftarrow Gr(G,k,d)$ setting $(1,v)$ into the sequence
@@ 5 Experimental results
This section reports the execution time and the size of the burning sequences returned by Algorithm 4 using Gr (Algorithm 3), GrP (Algorithm 6), and off-the-shelf optimization software (implementing Expressions (8) to (11)) over benchmark and synthetic graphs. All the real-world networks were obtained from the network repository [39] and the Stanford large network dataset collection [40]. Gr and GrP were implemented in C++ and compiled with GNU GCC. The ILP from Section 3 (Expressions (8) to (11)) was solved to optimality by Gurobi 11.0.0 [41] using default parameters and an aggressive presolve method. The execution times, even on small graphs, reported by Gurobi using the ILP from Section 2 (Expressions (1) to (6)) were very unpractical; thus, we did not report them. Table 1 shows all experiments’ hardware and software specifications. The implemented code can be consulted at https://github.com/jesgadiaz/GreedyBurning.
@@
Table 2 (small real-world graphs), Table 3 (medium and large real-world graphs), and Table 4 (square grids) show the execution time and the length of the solutions found by Gr, GrP, and Gurobi within a binary search (Algorithm 4). Column BFF reports the bounds used in the binary search; we used BFF algorithm to set $l=\lceil(len(s_{0})+2)/3\rceil$ and $h=len(s_{0})$, where $len(s_{0})$ is the length of the burning sequence returned by this algorithm; for reproducibility purposes, the first vertex selected by BFF is the one with the smallest label. Column BFS reports the time required to compute the distance matrix using breadth-first search (all reported times are in seconds). Using the ILP from Section 3 (Expressions (8) to (11)), Gurobi corroborated all the previously reported optimal solutions and found six new ones (tvshow, government, crocodile, DD6, DD349, and DD687). According to these results, the introduced ILP seems advantageous compared to the state-of-the-art ILP and CSPs [35]. For instance, Gurobi found the optimal burning sequence of tvshow in 473 seconds ($\sim$8 minutes) using the ILP from Section 3, while the same software could not find it in more than 48 hours using the state-of-the-art ILP and CSPs [35]. Besides, while the largest instance solved using the state-of-the-art formulations has 5908 vertices, Gurobi solved an instance of order 11631 in 1989 seconds ($\sim$33 minutes) using our formulation; we get an out-of-memory error for larger graphs. As expected, Gr was the fastest among all the tested algorithms. Its larger execution time was 1587 seconds ($\sim$26.5 minutes) over a graph of order 102400. A dash at Tables 2 to 4 indicates that the corresponding procedure was not executed because the memory requirements exceeded the hardware’s capacity or a time limit of 10 hours was estimated to be surpassed.
Table 2: Experimental results over small real-world networks. This table reports the execution time and quality of solutions found by Gr (Algorithm 3), GrP (Algorithm 6), and Gurobi (using Expressions (8) to (11)), all embedded into a binary search (Algorithm 4). Column BFF reports the time consumed by BFF algorithm and the bounds used in the binary search. Column BFS reports the time consumed by BFS to compute the distance matrix. Optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Graph BFF BFS Gurobi Gr GrP
name $n$ $m$ $b(G)$ $l$ $h$ $t$ $t$ $t$ size $t$ size $t$
karate 34 78 3 2 4 0.000 0.000 0.004 3 0.000 3 0.000
chesapeake 39 170 3 1 3 0.000 0.000 0.009 3 0.000 3 0.000
dolphins 62 159 4 2 6 0.000 0.000 0.036 4 0.000 4 0.010
rt-retweet 96 117 5 2 6 0.000 0.000 0.035 5 0.000 5 0.000
polbooks 105 441 4 2 5 0.000 0.000 0.047 4 0.000 4 0.015
adjnoun 112 425 4 2 5 0.000 0.000 0.044 4 0.004 4 0.015
ia-infect-hyper 113 2196 3 1 3 0.000 0.015 0.04 3 0.004 3 0.016
C125-9 125 6963 3 1 3 0.000 0.031 0.094 3 0.000 3 0.063
ia-enron-only 143 623 4 2 5 0.000 0.000 0.083 4 0.000 4 0.046
c-fat200-1 200 1534 7 3 7 0.000 0.015 0.235 7 0.000 7 0.220
c-fat200-2 200 3235 5 2 5 0.000 0.031 0.177 5 0.000 5 0.220
c-fat200-5 200 8473 3 1 3 0.000 0.062 0.079 3 0.004 3 0.095
sphere 258 1026 7 3 9 0.000 0.015 2.0 7 0.000 7 0.282
DD244 291 822 7 4 11 0.000 0.015 0.739 7 0.000 7 0.727
ca-netscience 379 914 6 3 8 0.000 0.031 0.632 7 0.003 6 0.644
infect-dublin 410 2765 5 2 6 0.000 0.062 0.614 5 0.015 5 1.2
c-fat500-1 500 4459 9 4 11 0.000 0.106 1.2 9 0.015 9 2.5
c-fat500-2 500 9139 7 3 8 0.000 0.203 1.0 7 0.015 7 3.0
c-fat500-5 500 23191 5 2 5 0.000 0.456 0.868 5 0.015 5 3.6
bio-diseasome 516 1188 7 5 13 0.000 0.046 1.6 7 0.031 7 2.0
web-polblogs 643 2280 5 3 8 0.000 0.109 2.2 6 0.046 5 6.0
DD687 725 2600 7 4 10 0.000 0.125 12 8 0.062 8 8.0
rt-twitter-copen 761 1029 7 3 9 0.000 0.093 4.7 7 0.078 7 9.3
DD68 775 2093 9 5 14 0.000 0.124 10 10 0.051 9 16
ia-crime-moreno 829 1475 7 3 8 0.000 0.112 12 7 0.078 7 20
DD199 841 1902 12 6 16 0.000 0.125 12 13 0.062 12 7.0
soc-wiki-Vote 889 2914 6 3 8 0.000 0.171 7.7 6 0.094 6 16
DD349 897 2087 12 6 18 0.000 0.157 12 13 0.062 12 20
DD497 903 2453 10 6 16 0.000 0.160 15 12 0.059 11 23
socfb-Reed98 962 18812 4 2 5 0.000 0.735 7.3 4 0.064 4 16
lattice3D 1000 2700 10 4 12 0.000 0.203 218 10 0.094 10 38
bal-bin-tree-9 1023 1022 10 4 10 0.000 0.125 3.1 10 0.031 10 5.8
delaunay-n10 1024 3056 9 4 11 0.000 0.236 33 10 0.078 9 27
stufe 1036 1868 12 5 15 0.000 0.176 122 12 0.078 12 22
lattice2D 1089 2112 13 7 19 0.000 0.187 172 14 0.078 13 40
bal-ter-tree-6 1093 1092 7 3 7 0.000 0.157 2.2 7 0.015 7 3.8
email-univ 1133 5451 5 2 6 0.000 0.359 13 5 0.094 5 23
econ-mahindas 1258 7513 5 2 6 0.000 0.488 54 5 0.109 5 31
ia-fb-messages 1266 6451 5 2 6 0.000 0.461 17 5 0.141 5 42
bio-yeast 1458 1948 9 4 11 0.000 0.343 29 9 0.234 9 143
Table 3: Experimental results over medium and large real-world networks. This table reports the execution time and quality of solutions found by Gr (Algorithm 3), GrP (Algorithm 6), and Gurobi (using Expressions (8) to (11)), all embedded into a binary search (Algorithm 4). Column BFF reports the time consumed by BFF algorithm and the bounds used in the binary search. Column BFS reports the time consumed by BFS to compute the distance matrix. Optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Graph BFF BFS Gurobi Gr GrP
name $n$ $m$ $b(G)$ $l$ $h$ $t$ $t$ $t$ size $t$ size $t$
tech-routers-rf 2113 6632 6 3 8 0.000 1.0 89 6 0.508 6 193
chameleon 2277 31421 6 3 8 0.000 3.4 61 6 0.659 6 350
tvshow 3892 17262 9 5 13 0.000 4.0 473 10 1.9 9 5201
facebook 4039 88234 4 2 5 0.000 14 29 4 0.533 4 369
DD6 4152 10320 16 9 25 0.000 3.4 1047 17 1.4 17 2868
squirrel 5201 198493 6 2 6 0.000 38 499 6 3.2 6 7394
politician 5908 41729 7 3 9 0.000 13 1239 7 3.7 7 8935
government 7057 89455 6 3 7 0.000 28 1233 6 4.2 6 10299
crocodile 11631 170918 6 3 7 0.000 84 1989 6 12 6 46798
athletes 13866 86858 ? 3 8 0.000 65 - 7 25 - -
company 14113 52310 ? 4 10 0.015 50 - 9 26 - -
musae-facebook 22470 171002 ? 4 11 0.015 208 - 8 72 - -
new-sites 27917 206259 ? 3 9 0.031 322 - 8 109 - -
deezer-europe 28281 92752 ? 4 12 0.031 208 - 10 105 - -
gemsec-deezer(RO) 41773 125826 ? 4 12 0.016 463 - 10 266 - -
gemsec-deezer(HU) 47538 222887 ? 4 10 0.046 780 - 8 376 - -
artist 50515 819306 ? 3 8 0.031 2035 - 6 418 - -
gemsec-deezer(HR) 54573 498202 ? 3 8 0.047 1631 - 7 426 - -
Table 4: Experimental results over square grids. This table reports the execution time and quality of solutions found by Gr (Algorithm 3), GrP (Algorithm 6), and Gurobi (using Expressions (8) to (11)), all embedded into a binary search (Algorithm 4). Column BFF reports the time consumed by BFF algorithm and the bounds used in the binary search. Column BFS reports the time consumed by BFS to compute the distance matrix. Optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Square grid BFF BFS Gurobi Gr GrP
name $n$ $m$ $b(G)$ $l$ $h$ $t$ $t$ $t$ size $t$ size $t$
10x10 100 180 6 3 8 0.000 0.001 0.161 7 0.015 6 0.031
20x20 400 760 10 5 13 0.000 0.027 5.2 11 0.017 10 0.897
30x30 900 1740 12 6 17 0.000 0.14 177 13 0.078 12 13
40x40 1600 3120 15 7 20 0.000 0.45 23742 16 0.191 15 92
50x50 2500 4900 ? 8 23 0.000 1.1 - 18 0.484 17 439
60x60 3600 7080 ? 9 26 0.011 2.2 - 21 1.1 19 1202
70x70 4900 9660 ? 10 29 0.000 4 - 22 1.5 21 3020
80x80 6400 12640 ? 11 31 0.000 7 - 25 3.2 24 5688
90x90 8100 16020 ? 12 34 0.000 11 - 28 5.1 25 12492
100x100 10000 19800 ? 12 36 0.013 17 - 28 5.8 27 32606
110x110 12100 23980 ? 13 38 0.013 25 - 30 10 - -
120x120 14400 28560 ? 14 41 0.015 36 - 32 15 - -
130x130 16900 33540 ? 15 43 0.015 50 - 34 20 - -
140x140 19600 38920 ? 16 46 0.046 69 - 35 27 - -
150x150 22500 44700 ? 16 48 0.046 93 - 37 35 - -
160x160 25600 50880 ? 17 51 0.062 123 - 39 44 - -
170x170 28900 57460 ? 18 53 0.063 161 - 41 56 - -
180x180 32400 64440 ? 19 55 0.063 208 - 42 69 - -
190x190 36100 71820 ? 19 57 0.094 259 - 44 85 - -
200x200 40000 79600 ? 20 60 0.141 322 - 45 123 - -
210x210 44100 87780 ? 20 60 0.126 397 - 47 124 - -
220x220 48400 96360 ? 21 63 0.157 487 - 48 150 - -
230x230 52900 105340 ? 22 64 0.181 591 - 50 215 - -
240x240 57600 114720 ? 23 67 0.188 713 - 51 211 - -
250x250 62500 124500 ? 23 67 0.345 838 - 52 300 - -
260x260 67600 134680 ? 23 69 0.219 1019 - 54 351 - -
270x270 72900 145260 ? 24 70 0.464 1174 - 55 402 - -
280x280 78400 156240 ? 25 73 0.456 1388 - 57 471 - -
290x290 84100 167620 ? 25 73 5.2 1627 - 58 641 - -
300x300 90000 179400 ? 25 74 19 1954 - 59 983 - -
310x310 96100 191580 ? 26 76 53 2306 - 60 1522 - -
320x320 102400 204160 ? 26 78 35 2687 - 62 1587 - -

### v3

@@ 4.2 A greedy heuristic for GBP
It is well known that $b(P_{n})=\lceil n^{1/2}\rceil$ for any path $P_{n}$ of order $n$ [1]. So, given a path $P_{n}$, we can run Gr with $p=b(P_{n})$, which is equivalent to execute Algorithm 4 with $l=h=b(P_{n})$. For the rest of the proof, let us consider paths $P_{n}$ with the largest possible order for any given burning number, namely, paths $P_{n}$ with $n=z^{2}$ for any $z\in\mathbb{Z}^{+}$. Also, observe that the maximum number of vertices a vertex $v$ can cover in a path is $|N_{r}[v]|=2r+1$, for any given $r$.
The proof is divided into three parts. The first considers the initial iteration of the for loop of Gr ($r=p-1$); the second considers iterations $r=p-2$ to $r=2$, and the last one considers iterations $r=1$ and $r=0$. In the first iteration, $r=p-1$, Gr selects a vertex that covers exactly $12r+$ 1 vertices because the path’s length, $(r+1)^{2}$, is bounded by $12r+$ 1, i.e.,
@@
At iterations $r=p-2$ to $r=2$, Gr selects a vertex with maximum coverage from the largest subpath induced by removing the covered vertices. The largest subpath has at least $12r+$ 1 vertices, i.e.,
@@
where $(r+2)^{2}$ is the number of uncovered vertices two iterations before, and $12(r+1)+$ 1 is the number of newly covered vertices at the previous iteration. Thus, the numerator is the number of uncovered vertices at the beginning of iteration $r$. This way, at iteration $r$, Gr selects a vertex that covers $12r+$ 1 vertices, none of which were covered in the previous iterations. Notice that ties can be broken at random. By mathematical induction, Eq. (40) holds for any integer $r\geq 2$.
At iteration $r=1$, the number of previously covered vertices equals $\sum_{r=2}^{p-1}(2r+1)=p^{2}-4$. Namely, at this point, only four vertices are left uncovered. By direct inspection, there will be cases where only two or three new vertices can be covered at iterations $r=1$ and $r=0$, leaving the rest uncovered (See Figure 6). Therefore, the constructed solution might not be a burning sequence.
Figure 6: A non-burning sequence returned by Gr over $P_{25}$ with $p=b(P_{25})=5$. At the end, two vertices are uncovered.
@@
Since a cycle $C_{n}$ is a path $P_{n}$ with its endpoints joined, it follows that $b(C_{n})=\lceil n^{1/2}\rceil$ [1]. So, we can run Gr with $p=b(C_{n})$, which is equivalent to execute Algorithm 4 with $l=h=b(C_{n})$. The first vertex selected by Gr maximizes the coverage. Then, removing the covered vertices induces a path $P$. By Proposition 4, Gr might return a non-burning sequence of length $b(C_{n})-1$ over $P$, which is concatenated to the first selected vertex, becoming a non-burning sequence of length $b(C_{n})$ for $C_{n}$. ∎
By Proposition 4 and Corollary 1, Gr might not seem very appealing. However, as reported in Section 5, it performed unexpectedly well over benchmark and synthetic graphs. Anyhow, we can do better by repeating Gr up to $n$ times and selecting a different initial vertex every time; we refer to this heuristic as GrP (Algorithm 6). Once the first burning sequence of length $p$ is found (line 3), exploring the remaining vertices is unnecessary. Nevertheless, the complexity of GrP embedded into a binary search (Algorithm 4) is $\mathcal{O}(n^{4}\log n)$, which is impractical for sufficiently large graphs.
@@
1:A graph $G=(V,E)$, its distance matrix $D$, and a guess $p$ on $b(G)$
@@
4:     $S\leftarrow Gr(G,p,d)$ with $v$ as the first selected vertex
@@ 5 Experimental results
This section reports the execution time and the size of the burning sequences returned by Algorithm 4 using Gr (Algorithm 3) and GrP (Algorithm 6) over benchmark and synthetic graphs. Off-the-shelf optimization software implementing ILP-COV (Expressions (11) to (17)) was used to compute optimal solutions; the upper bound $U$ being equal to the length of the best feasible solution returned by Gr and GrP. All the real-world networks were obtained from the network repository [22] and the Stanford large network dataset collection [23]. Gr and GrP were implemented in C++ and compiled with GNU GCC. ILP-COV was solved to optimality by Gurobi 11.0.0 [24] using default parameters. The execution times of Gurobi using ILP-PROP and ILP-CMCP (Algorithm 1) were larger; thus, we did not report them. Table 1 shows all experiments’ hardware and software specifications. The implemented codes can be consulted at https://github.com/jesgadiaz/GreedyBurning.
@@
Table 2 (real-world graphs with less than 6000 vertices), Table 3 (real-world graphs with order between 7057 and 88784), and Table 4 (square grids) show the execution time and the length of the solutions found by Gr and GrP within a binary search (Algorithm 4). Column BFF reports the bounds used in the binary search; we used BFF algorithm to set $l=\lceil(s(S_{0})+2)/3\rceil$ and $h=s(S_{0})-1$, where $s(S_{0})$ is the size of the burning sequence returned by this algorithm; for reproducibility purposes, the first vertex selected by BFF is the one with the smallest label. Column BFS reports the time required to compute the distance matrix using breadth-first search (all reported times are in seconds). Column Gurobi reports the burning number computed by this solver using ILP-COV, where the upper bound $U$ is the size of the best feasible solution found by Gr and GrP.
Using ILP-COV, Gurobi corroborated all the previously reported optimal solutions and found six new ones (tvshow, government, crocodile, DD6, DD349, and DD687). According to these results, ILP-COV seems advantageous compared to the state-of-the-art ILP and CSPs [17]. For instance, Gurobi found the optimal burning sequence of tvshow in 62 seconds using ILP-COV, while the same software could not find it in more than 48 hours using the ILP and CSPs reported by [17]. Besides, while the largest instance solved by the aforementioned models has 5908 vertices, Gurobi solved an instance of order 11631 in 841 seconds ($\sim$14 minutes) using ILP-COV; we get an out-of-memory error for larger graphs. For the graphs with a dash symbol at column $t$ of the Gurobi column, we report the burning number found by a recent state-of-the-art exact algorithm [25]; although this algorithm is faster than Gurobi using ILP-COV, we are reporting the running time of the latter for comparison and corroboration purposes. Regarding Gr and GrP, the former was the fastest among all the tested algorithms, requiring 3282 seconds ($\sim$55 minutes) over a graph of order 102400. A dash at Tables 2 to 4 indicates that the corresponding procedure was not executed because the memory requirements exceeded the hardware’s capacity or a time limit of 10 hours was estimated to be surpassed. In all tables, optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Table 2: Experimental results over real-world networks with less than 6000 vertices.
Graph Gurobi BFF BFS Gr GrP
name $n$ $m$ $b(G)$ $t$ $l$ $s(S_{0})$ $t$ $t$ size $t$ size $t$
karate 34 78 3 0.015 2 4 0.000 0.000 3 0.000 3 0.000
chesapeake 39 170 3 0.000 1 3 0.000 0.000 3 0.000 3 0.001
dolphins 62 159 4 0.000 2 6 0.000 0.000 4 0.001 4 0.004
rt-retweet 96 117 5 0.016 2 6 0.000 0.000 5 0.001 5 0.024
polbooks 105 441 4 0.016 2 5 0.000 0.000 4 0.000 4 0.018
adjnoun 112 425 4 0.016 2 5 0.000 0.000 4 0.000 4 0.032
ia-infect-hyper 113 2196 3 0.032 1 3 0.000 0.015 3 0.000 3 0.018
C125-9 125 6963 3 0.125 1 3 0.000 0.031 3 0.000 3 0.055
ia-enron-only 143 623 4 0.031 2 5 0.000 0.000 4 0.002 4 0.046
c-fat200-1 200 1534 7 0.044 3 7 0.000 0.015 7 0.002 7 0.307
c-fat200-2 200 3235 5 0.029 2 5 0.000 0.031 5 0.003 5 0.240
c-fat200-5 200 8473 3 0.016 1 3 0.000 0.062 3 0.000 3 0.105
sphere 258 1026 7 1.6 3 9 0.000 0.015 7 0.006 7 0.488
DD244 291 822 7 0.244 4 11 0.000 0.015 7 0.007 7 0.756
ca-netscience 379 914 6 0.077 3 8 0.000 0.031 7 0.016 6 0.643
infect-dublin 410 2765 5 0.204 2 6 0.000 0.062 5 0.018 5 1.9
c-fat500-1 500 4459 9 0.174 4 11 0.000 0.106 9 0.015 9 2.7
c-fat500-2 500 9139 7 0.142 3 8 0.000 0.203 7 0.018 7 3.2
c-fat500-5 500 23191 5 0.255 2 5 0.000 0.456 5 0.014 5 3.9
bio-diseasome 516 1188 7 0.174 5 13 0.000 0.046 7 0.028 7 2.2
web-polblogs 643 2280 5 0.440 3 8 0.000 0.109 6 0.049 5 5.0
DD687 725 2600 7 4.1 4 10 0.000 0.125 8 0.053 8 14
rt-twitter-copen 761 1029 7 0.852 3 9 0.000 0.093 7 0.061 7 14
DD68 775 2093 9 2.1 5 14 0.000 0.124 10 0.059 9 16
ia-crime-moreno 829 1475 7 4.2 3 8 0.000 0.112 7 0.075 7 20
DD199 841 1902 12 1.7 6 16 0.000 0.125 13 0.066 12 13
soc-wiki-Vote 889 2914 6 1.4 3 8 0.000 0.171 6 0.065 6 16
DD349 897 2087 12 1.9 6 18 0.000 0.157 13 0.076 12 8.6
DD497 903 2453 10 2.1 6 16 0.000 0.160 12 0.086 11 9.4
socfb-Reed98 962 18812 4 2.4 2 5 0.000 0.735 4 0.077 4 16
lattice3D 1000 2700 10 116 4 12 0.000 0.203 10 0.089 10 34
bal-bin-tree-9 1023 1022 10 0.883 4 10 0.000 0.125 10 0.024 10 8.2
delaunay-n10 1024 3056 9 5.6 4 11 0.000 0.236 10 0.072 9 27
stufe 1036 1868 12 13 5 15 0.000 0.176 12 0.076 12 30
lattice2D 1089 2112 13 22 7 19 0.000 0.187 14 0.099 13 20
bal-ter-tree-6 1093 1092 7 0.343 3 7 0.000 0.157 7 0.022 7 4.7
email-univ 1133 5451 5 3.1 2 6 0.000 0.359 5 0.105 5 29
econ-mahindas 1258 7513 5 3.5 2 6 0.000 0.488 5 0.126 5 40
ia-fb-messages 1266 6451 5 3.7 2 6 0.000 0.461 5 0.152 5 54
bio-yeast 1458 1948 9 7.4 4 11 0.000 0.343 9 0.237 9 141
tech-routers-rf 2113 6632 6 9.4 3 8 0.000 1.0 6 0.314 6 194
chameleon 2277 31421 6 14.1 3 8 0.000 3.4 6 0.444 6 334
tvshow 3892 17262 9 62 5 13 0.000 4.0 10 1.9 9 3057
facebook 4039 88234 4 9.2 2 5 0.000 14 4 0.547 4 377
DD6 4152 10320 16 72 9 25 0.000 3.4 17 1.3 17 821
squirrel 5201 198493 6 124 2 6 0.000 38 6 2.6 6 8853
politician 5908 41729 7 113 3 9 0.000 13 7 3.8 7 9068
Table 3: Experimental results over real-world networks with order between 7057 and 88784.
Graph Gurobi BFF BFS Gr GrP
name $n$ $m$ $b(G)$ $t$ $l$ $s(S_{0})$ $t$ $t$ size $t$ size $t$
government 7057 89455 6 178 3 7 0.000 28 6 5.3 6 14973
crocodile 11631 170918 6 841 3 7 0.000 84 6 16 6 81669
athletes 13866 86858 7 - 3 8 0.000 65 7 23 - -
company 14113 52310 9 - 4 10 0.015 50 9 23 - -
musae-facebook 22470 171002 8 - 4 11 0.015 208 8 68 - -
new-sites 27917 206259 8 - 3 9 0.031 322 8 95 - -
deezer-europe 28281 92752 10 - 4 12 0.031 208 10 93 - -
gemsec-deezer(RO) 41773 125826 10 - 4 12 0.016 463 10 241 - -
gemsec-deezer(HU) 47538 222887 8 - 4 10 0.046 780 8 325 - -
artist 50515 819306 6 - 3 8 0.031 2035 6 274 - -
gemsec-deezer(HR) 54573 498202 7 - 3 8 0.047 1631 7 433 - -
soc-brightkite 56739 212945 9 - 3 9 0.062 963 9 437 - -
socfb-OR 63392 816886 8 - 3 9 0.051 2616 8 795 - -
soc-slashdot 70068 358647 7 - 3 8 0.031 1718 7 1401 - -
soc-BlogCatalog 88784 2093195 5 - 2 6 3.9 8600 5 3315 - -
Table 4: Experimental results over square grids.
Graph Gurobi BFF BFS Gr GrP
name $n$ $m$ $b(G)$ $t$ $l$ $s(S_{0})$ $t$ $t$ size $t$ size $t$
10x10 100 180 6 0.099 3 8 0.000 0.001 7 0.001 6 0.023
20x20 400 760 10 1.6 5 13 0.000 0.027 11 0.018 10 1.7
30x30 900 1740 12 46 6 17 0.000 0.14 13 0.079 12 13
40x40 1600 3120 15 58 7 20 0.000 0.45 16 0.202 15 93
50x50 2500 4900 17 302 8 23 0.000 1.1 18 0.469 17 415
60x60 3600 7080 19 1107 9 26 0.011 2.2 21 0.873 19 1219
70x70 4900 9660 21 20073 10 29 0.000 4 22 1.5 21 3068
80x80 6400 12640 ? - 11 31 0.000 7 25 3.2 24 8241
90x90 8100 16020 ? - 12 34 0.000 11 28 5.0 25 1756
100x100 10000 19800 ? - 12 36 0.013 17 28 7.4 27 22301
110x110 12100 23980 ? - 13 38 0.013 25 30 10 - -
120x120 14400 28560 ? - 14 41 0.015 36 32 15 - -
130x130 16900 33540 ? - 15 43 0.015 50 34 20 - -
140x140 19600 38920 ? - 16 46 0.046 69 35 27 - -
150x150 22500 44700 ? - 16 48 0.046 93 37 35 - -
160x160 25600 50880 ? - 17 51 0.062 123 39 44 - -
170x170 28900 57460 ? - 18 53 0.063 161 41 56 - -
180x180 32400 64440 ? - 19 55 0.063 208 42 70 - -
190x190 36100 71820 ? - 19 57 0.094 259 44 86 - -
200x200 40000 79600 ? - 20 60 0.141 322 45 105 - -
210x210 44100 87780 ? - 20 60 0.126 397 47 126 - -
220x220 48400 96360 ? - 21 63 0.157 487 48 182 - -
230x230 52900 105340 ? - 22 64 0.181 591 50 218 - -
240x240 57600 114720 ? - 23 67 0.188 713 51 257 - -
250x250 62500 124500 ? - 23 67 0.345 838 52 301 - -
260x260 67600 134680 ? - 23 69 0.219 1019 54 353 - -
270x270 72900 145260 ? - 24 70 0.464 1174 55 404 - -
280x280 78400 156240 ? - 25 73 0.456 1388 57 478 - -
290x290 84100 167620 ? - 25 73 5.2 1627 58 575 - -
300x300 90000 179400 ? - 25 74 18.6 1954 59 1010 - -
310x310 96100 191580 ? - 26 76 53.4 2306 60 2101 - -
320x320 102400 204160 ? - 26 78 35 2687 62 3282 - -